OpenStudy (anonymous):

Please help me. I suck at math >.< Find the zeros of f(x) = x2 - 9 Find the zeros of the function h(x) = 3x2 + 13x + 4 Find the zeros of the function g(x) = x2 - 6x - 16 Find the zeros of f(x) = x2 + 4x - 5

OpenStudy (anonymous):

equate f(x) , g(x) , h(x) to zero and find

OpenStudy (anonymous):

for example i'll solve first one for u...... x^2 - 9=0 X^2=9 x=9^(1/2) x= +3, -3

OpenStudy (anonymous):

But could you explain it to me? I dont really get it.

OpenStudy (anonymous):

first we hv x^2 = 9 Get it???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Then take square root on both sides Get it??

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so sqaure root of x^2 is x and square root of 9 are +3 and -3 Now???

OpenStudy (anonymous):

Yeah okay thank you! Okay, so can you explain the second one? Theres an extra step in that one. right?

OpenStudy (anonymous):

ya in that one u got2 create factors do u know the method of splitting the middle term????

OpenStudy (anonymous):

eh, I don't really get it.

OpenStudy (anonymous):

see u got2 split the middle term i.e. 13x into two terms so that sum of the terms is 13x and product is 3*4=12 compare the given equation with std. one 3x^2 + 13x + 4 (ax^2 + bx + c) so b= m + n and a*c=m*n Did u get it???

OpenStudy (anonymous):

Yeah I get that

OpenStudy (anonymous):

so will u try it by urself now????

OpenStudy (anonymous):

coz the rest 2 r also of same type

OpenStudy (anonymous):

tell will u try or shall i wait???

OpenStudy (anonymous):

Yeah ill try, thank you for the help!(:

OpenStudy (anonymous):

ur welcum

OpenStudy (s3a):

Just in case you're going about this blindly, "finding zeroes" means finding all the points on the function that lie somewhere along the x-axis aka at some x-value with y = 0. You could also use the quadratic formula if you want to remove the thinking and just keep this "plug-and-play".

OpenStudy (anonymous):

Oh, thank you :)

OpenStudy (s3a):

No problem. :)